# ArtNeRF: A Stylized Neural Field for 3D-Aware Cartoonized Face Synthesis

### v1

@@ 3.2 Self-supervised Style Encoder
The overall structure of the style encoder is illustrated in Fig. 3. A pretrained VGG19 is utilized to extract style features $\boldsymbol{f_{s}}$ from input images, followed by a compress-CNN to reduce the dimension of $\boldsymbol{f_{s}}$ to 512. The 512-dim vector serves as the style code $\boldsymbol{z_{s}}$. To facilitate the contrastive learning process, a projection head is applied to further map $\boldsymbol{z_{s}}$ to its representation vectors. During training, each batch contains $2N$ images, where $\boldsymbol{X_{i}},\boldsymbol{X_{j}}$ are positive samples ($\boldsymbol{X_{i}}$ is a style image and $\boldsymbol{X_{j}}$ is an augmented sample via affine transformation), and the remaining $12N-$ 1 images serve as negative samples. We use the following objective function to optimize the compress-CNN :
$\ell_{i,j}^{CL}=-\log\frac{{\exp({\mathop{\rm sim}\nolimits}({{\boldsymbol{z_{i}}}},{{\boldsymbol{z_{j}}}})/\tau)}}{{\sum\limits_{k=1,k\neq i,k\neq j}^{2N}{\exp({\mathop{\rm sim}\nolimits}({{\boldsymbol{z_{i}}}},{{\boldsymbol{z_{k}}}})/\tau)}}}$ (2)
where $\rm sim(\cdot,\cdot)$ is the cosine similarity between embeddings, $\tau$ is the temperature coefficient, and $(\boldsymbol{z_{i}},\boldsymbol{z_{j}})$ represents the embeddings for $(\boldsymbol{X_{i}},\boldsymbol{X_{j}})$. After training, augmented samples of the same style image will have style codes rich in style semantics but devoid of spatial structure, since they are pushed closer in the embedding space and their original spatial features are neutralized.
@@ 3.3 Conditional Generative Radiance Field
$\begin{gathered}{\boldsymbol{W_{fused}}=\rm{Concat}(\boldsymbol{W_{bb}};\boldsymbol{W_{nr}}})\\ {\boldsymbol{W_{bb}}}={\boldsymbol{\alpha}[:k]}\odot{{\boldsymbol{w_{s}}[:k,:]}}+(1-{\boldsymbol{\alpha}[:k]})\odot{{\boldsymbol{w_{f}}[:k,:]}}\\ {\boldsymbol{W_{nr}}}={\boldsymbol{\alpha}[k:]}\odot\rm{Trans}({{\boldsymbol{w_{s}}[k:,:]}})+(1-{\boldsymbol{\alpha}[k:]})\odot\rm{Trans}({{\boldsymbol{w_{f}}[k:,:]}})\end{gathered}$ (3)
[…]
Conditional Radiance Field with Dense Skip Connections. The proposed conditional radiance field takes as input not only 3D position $\boldsymbol{x}$ and viewing direction $\boldsymbol{v}$ but also a fused conditioning latent code $\boldsymbol{W_{bb}}$. Therefore, properly injecting $\boldsymbol{W_{bb}}$ into the backbone is crucial for the generation performance. As is shown in Fig.2, the backbone consists of two parts: a sequence of $n$ FiLM layers and dense skip connections. The FiLM layer sequence can be formulated as follows:
@@
where $\boldsymbol{x_{i}}$ is the input of the $i$-th FiLM layer, $\boldsymbol{W_{i}},\boldsymbol{b_{i}}$ are learnable parameters and $\gamma_{i},\beta_{i}$ are modulation coefficients projected from $\boldsymbol{W_{bb}}$. Note that $\phi_{n-1}$ also takes $\boldsymbol{v}$ as input to model view-dependant appearance, we omit it for brevity. To mitigate the ripple-like artifacts observed during training pi-GAN, we draw inspiration from StyleGAN2’s [12] improvements to StyleGAN. We discard the progressive growing training strategy used in pi-GAN and optimize the generator with a structure featuring dense skip connections. This modification ensures that feature maps from different layers can mutually contribute to the final output, thereby increasing the strength of gradient back-propagation and preventing training collapse. Optimized formulas for volume density and feature calculation is shown in Eq.5:
@@ 3.4 Neural Rendering Module
The integration of the neural rendering module can remarkably enhance the expressiveness of the generator, enabling the synthesis of high-quality images at higher resolutions with faster inference speed. GIRAFFE [21] first introduces a neural renderer composed of 2D CNNs into the model. However, EG3D [2] indicates that overly deep 2D CNNs and excessive $3\times 3$ convolution operations undermine the 3D consistency of the final results. Therefore, our neural rendering module is composed of shallow $1\times 1$ convolutions. $1\times 1$ convolutions enhance the network’s ability to model information between channels of 2D feature maps and help avoid the fusion of local spatial contexture in feature maps, further ensuring multi-view consistency. The proposed neural rendering module leverages ModConv and upsampling layers in StyleGAN2, with a to_rgb layer to facilitate the reuse of features between adjacent blocks. Introducing ModConv allows for the reuse of $\boldsymbol{W_{nr}}$ during super-resolution, which can refine the details of low-resolution results. Given a 2D low-resolution feature map $\boldsymbol{F}\in\mathbb{R}^{{M_{f}}\times 32\times 32}$, we can generate the final synthesized image $\boldsymbol{X_{syn}}\in\mathbb{R}^{3\times 128\times 128}$ following Fig. 4, where the $A$ blocks are affine transformations applied to inject $\boldsymbol{W_{nr}}$ into the neural rendering module and the $h_{rgb}$ is a clamp operation.

### v2

@@ 3.2 Self-supervised Style Encoder
The overall structure of the style encoder is illustrated in Fig. 3. A pretrained VGG19 is utilized to extract style features $\boldsymbol{f_{s}}$ from input images, followed by a compress-CNN to reduce the dimension of $\boldsymbol{f_{s}}$ to 512. The 512-dim vectors serve as the style codes $\boldsymbol{z_{s}}$. To facilitate the contrastive learning process, a projection head is applied to further map $\boldsymbol{z_{s}}$ to their representation vectors $\boldsymbol{u_{s}}$. During training, each batch contains $2N$ images, where $\boldsymbol{X_{i}},\boldsymbol{X_{j}}$ are positive samples ($\boldsymbol{X_{i}}$ is a style image and $\boldsymbol{X_{j}}$ is an augmented sample via affine transformation), and the remaining $22N-$ 2 images serve as negative samples. We use the following objective function to optimize the compress-CNN :
$\ell_{i,j}^{CL}=-\log\frac{{\exp({\mathop{\rm sim}\nolimits}({{\boldsymbol{u_{i}}}},{{\boldsymbol{u_{j}}}})/\tau)}}{{\sum\limits_{k=1,k\neq i,k\neq j}^{2N}{\exp({\mathop{\rm sim}\nolimits}({{\boldsymbol{u_{i}}}},{{\boldsymbol{u_{k}}}})/\tau)}}}$ (2)
where $\rm sim(\cdot,\cdot)$ is the cosine similarity between embeddings, $\tau$ is the temperature coefficient, and $(\boldsymbol{u_{i}},\boldsymbol{u_{j}})$ represents the contrastive learning representations for $(\boldsymbol{X_{i}},\boldsymbol{X_{j}})$. After training, augmented samples of the same style image will have style codes rich in style semantics but devoid of spatial structure, since they are pushed closer in the embedding space and their original spatial features are neutralized.
@@ 3.3 Conditional Generative Radiance Field
$\begin{gathered}{\boldsymbol{W_{fused}}=\rm{Concat}(\boldsymbol{W_{bb}};\boldsymbol{W_{nr}}})\\ {\boldsymbol{W_{bb}}}={\boldsymbol{\alpha}[:k]}\odot{{\boldsymbol{w_{s}}[:k,:]}}+(1-{\boldsymbol{\alpha}[:k]})\odot{{\boldsymbol{w_{f}}[:k,:]}}\\ {\boldsymbol{W_{nr}}}={\boldsymbol{\alpha}[k:]}\odot\rm{Trans}({{\boldsymbol{w_{s}}[\mathnormal{k}:,:]}})+(1-{\boldsymbol{\alpha}[\mathnormal{k}:]})\odot\rm{Trans}({{\boldsymbol{w_{f}}[\mathnormal{k}:,:]}})\end{gathered}$ (3)
where $bb$ and $nr$ denote backbone and neural rendering, $\rm{Concat}(\cdot;\cdot)$ and $\odot$ denote channel-wise concatenation and element-wise multiplication, [:]delimited-[]:[:][ : ] represent the slicing operation in PyTorch. In practice, $k$ is the number of layers in the backbone, $\boldsymbol{\alpha}$ is a learnable weight vector with a shape of $[n]$, $\boldsymbol{w_{f}}$ and $\boldsymbol{w_{s}}$ both have a shape of $[n,256]$, where $n$ is the number of layers requiring style mixing. To flexibly adjust the degree of stylization, we introduce a split index $i$ in SBM. When $i$ is specified, $\boldsymbol{\alpha}[:i]=1$ and $\boldsymbol{\alpha}[i:]$ remains unchanged. This strategy ensures layers with indices less than $i$ only affected by $\boldsymbol{w_{f}}$, while layers with indices greater than $i$ influenced by both $\boldsymbol{w_{f}}$ and $\boldsymbol{w_{s}}$. Consequently, we can perform style blending on the backbone and the neural rendering module (Sec 3.4). Note that directly injecting style vector into the neural rendering module may be improper as its feature space differs from the backbone. Hence, we apply projection units to $\boldsymbol{w_{f}}$ and $\boldsymbol{w_{s}}$ to refine them before the injection operation.
Conditional Radiance Field with Dense Skip Connections. The proposed conditional radiance field takes as input not only 3D positions $\boldsymbol{x}$ in the camera coordinate system and a camera pose $\boldsymbol{p}$ but also a fused conditioning latent code $\boldsymbol{W_{bb}}$. Therefore, properly injecting $\boldsymbol{W_{bb}}$ into the backbone is crucial for the generation performance. As is shown in Fig.2, the backbone consists of two parts: a sequence of $n$ FiLM layers and dense skip connections. The FiLM layer sequence can be formulated as follows:
@@
where $\boldsymbol{x_{i}}$ is the input of the $i$-th FiLM layer, $\boldsymbol{W_{i}},\boldsymbol{b_{i}}$ are learnable parameters and $\gamma_{i},\beta_{i}$ are modulation coefficients projected from $\boldsymbol{W_{bb}}$. Note that $\phi_{n-1}$ also takes the viewing direction $\boldsymbol{v}$ as input to model view-dependant appearance, we omit it for brevity. To mitigate the ripple-like artifacts observed during training pi-GAN, we draw inspiration from StyleGAN2’s [12] improvements to StyleGAN. We discard the progressive growing training strategy used in pi-GAN and optimize the generator with a structure featuring dense skip connections. This modification ensures that feature maps from different layers can mutually contribute to the final output, thereby increasing the strength of gradient back-propagation and preventing training collapse. Optimized formulas for volume density and feature calculation is shown in Eq.5:
@@ 3.4 Neural Rendering Module
The integration of the neural rendering module can remarkably enhance the expressiveness of the generator, enabling the synthesis of high-quality images at higher resolutions with faster inference speed. GIRAFFE [21] first introduces a neural renderer composed of 2D CNNs into the model. However, EG3D [2] indicates that overly deep 2D CNNs and excessive $3\times 3$ convolution operations undermine the 3D consistency of the final results. Therefore, our neural rendering module is composed of shallow $1\times 1$ convolutions. $1\times 1$ convolutions enhance the network’s ability to model information between channels of 2D feature maps and help avoid the fusion of local spatial contexture in feature maps, further ensuring multi-view consistency. The proposed neural rendering module leverages ModConv and upsampling layers in StyleGAN2, with a to_rgb layer to facilitate the reuse of features between adjacent blocks. Introducing ModConv allows for the reuse of $\boldsymbol{W_{nr}}$ during super-resolution, which can refine the details of low-resolution results. Given a 2D low-resolution feature map $\boldsymbol{F}\in\mathbb{R}^{{M_{f}}\times 32\times 32}$, we can generate the final synthesized image $\boldsymbol{X_{syn}}\in\mathbb{R}^{3\times 128\times 128}$ following Fig. 4, where the $A$ blocks are affine transformations applied to inject $\boldsymbol{W_{nr}}$ into the neural rendering module, $\mathbf{w_{0}}\dots\mathbf{w_{5}}$ are weight parameters for $1\times 1$ convolutions and the $h_{rgb}$ is a clamp operation.